A father is 6 times as old as his son and 3 times as old as his daughter. If the daughter is 5 years older than the son ,how old is each?
Answer by palanisamy(496)

Let the age of the son = x
Then the age of the daughter = x+5
Given,A father is 6 times as old as his son and 3 times as old as his daughter.
6x = 3(x+5)
6x = 3x+15
6x-3x = 15
3x = 15
x = 15/3
x = 5
Therefore,the age of the son = 5 years
The age of the daughter = x+5 = 5+5 = 10 years
the age of the son = x
The age of the father = 6x = 6*5 = 30 years